On 2 February, 2014, 1 Swiss Franc (CHF) was trading for USD1.0044 in the spot market. On the same day of 2019, CHF1 = USD1.1031. What happened to CHF against USD? What happened to USD against CHF? What is the annual percentage change in CHF? In USD?
Answer #1

In 2014, 1 CHF = 1.0044 USD

in 2019, 1 CHF = 1.1031 USD

Hence, CHF has appreciated against USD

USD has depreciated agains CHF

% Change in USD = ((1.1031 - 1/1.0044)/(1.0044)) =0.1070

Annual % change = (1+0.1070)1/5 -1 = 0.0205 or 2.05%

% Change in USD = ((1/1.1031 - 1/1.0044)/(1/1.0044)) =-0.0895

Annual % change = (1-0.0894)1/5 -1 = -0.0186 or -1.86%